A person uses 770 kcal on a long hike. Calculate the energy used for the hike in each of the following energy units.

What is the energy in joules?

Express the energy in joules to two significant figures.

What is the energy in kilojoules?

Express the energy in kilojoules to two significant figures.
Answer #1

1 kcal= 4184 Joules

770 kcal= 770×4184 = 3,221,680 Joules i.e.3.22×(10)^6

3,200,000 Joules ( upto two significant figures)

1kcal = 4.184 kilojoules

770 kcal= 770×4.184 = 3221.68 kilojoules i.e.3.221×(10)^3

3200 kilojoules ( upto two significant figures)